OpenStudy (anonymous):

What is the area of a circle with diameter 12 inches? 144piin2 6piin2 12piin2 36piin2

OpenStudy (anonymous):

who ever get this correct get a medal

OpenStudy (tylerd):

radius squared times pi

OpenStudy (imnotgoodatmath):

wierd? I got 113.1 in^2

OpenStudy (anonymous):

congrats you get a medal

OpenStudy (imnotgoodatmath):

dangit :D

OpenStudy (anonymous):

A=pi*r^2 =3.14*6^2 =3.14* 36 =113.04 sq. inches

OpenStudy (anonymous):

so its 144 @lupita1995

OpenStudy (anonymous):

i say so

OpenStudy (anonymous):

A

OpenStudy (anonymous):

well congrats

OpenStudy (anonymous):

@lupita1995 you get a medal

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

i don't really know what to choose A or C

OpenStudy (anonymous):

it is A

OpenStudy (tylerd):

D

OpenStudy (tylerd):

diameter is 12 so the radius is 6 6^2 = 36 36pi

OpenStudy (tylerd):

sorry i read it wrong